# Masking Augmentation for Supervised Learning

### v2

@@ Abstract
Pre-training using random masking has emerged as a novel trend in training techniques. However, supervised learning faces a challenge in adopting masking augmentations, primarily due to unstable training. In this paper, we propose a novel way to involve masking augmentations dubbed Masked Sub-model (MaskSub). MaskSub consists of the main-model and sub-model; while the former enjoys conventional training recipes, the latter leverages the benefit of strong masking augmentations in training. MaskSub addresses the challenge by mitigating adverse effects through a relaxed loss function similar to a self-distillation loss. Our analysis shows that MaskSub improves performance, with the training loss converging even faster than regular training, which suggests our method facilitates training. We further validate MaskSub across diverse training recipes and models, including DeiT-III, MAE fine-tuning, CLIP fine-tuning, ResNet, and Swin Transformer. Our results show that MaskSub consistently provides significant performance gains across all the cases. MaskSub provides a practical and effective solution for introducing additional regularization under various training recipes. Code available at https://github.com/naver-ai/augsub
@@ 1 Introduction
Supervised learning is the most basic and effective way to train a network to achieve high performance on a target task. To improve supervised learning, diverse regularizations are developed and used as training recipes [45, 47, 54], which represent a group of sophisticatedly tuned regularizations to maximize learning performance. Supervised learning has always held an advantage over self-supervised learning [7, 4] based on the benefit of supervision. However, emergence of Vision Transformer (ViT) [15] and Masked Image Modeling (MIM) [1, 21, 38] is changing this trends. ViT, which lacks inductive bias than convolution networks, poses a lot of challenges to generalization performance for supervised learning. On the other hand, MIMs such as MAE [21] rise as an alternative pre-training method for ViT by achieving competitive performance with supervised learning recipes. Although a recent study [47] shows that new supervised learning outperforms MIMs, the gap is not significant. Thus, MIMs are still a strong competitor of supervised learning methods.
@@
MIM masks random areas of an input image and forces the network to infer the masked area using the remaining area. A representative characteristic of MIM is high mask ratios over 50%. Although MIM also works at small mask ratios, it shows remarkable performance when trained with a high mask ratio. The high mask ratio is a major difference between MIM from supervised learning, since this high mask ratio, over 50%, is not allowed in supervised learning. Supervised learning also has utilized random masking as an augmentation [17, 66] but it significantly degrades performance when the masking ratio is high. In other words, supervised learning is not applicable for strong masking augmentation. We conjecture that it is a major problem of the current supervised learning recipe and there is room for improvement by enabling strong masking augmentation for supervised learning.
Our goal is to improve supervised learning with strong mask augmentation over 50% while maintaining the stability of training recipes. To this end, we introduce a training framework using a “sub-model” alongside the main-model; throughout this paper, we use the term “sub-model” to describe a partial model based on the ”main-model”, in which some network features do not engage in training. The main-model uses standard training recipes [54, 47]. On the other hand, the sub-model utilizes mask augmentation. We name our method as Masked Sub-model (MaskSub). In the example of Figure 1, 50% random masking is applied to the sub-model as done in MAE [21]. As in Figure 1 (b), applying the 50% random masking to the main-model may lead to degraded performance. In contrast, as in Figure 1 (c), MaskSub utilizes the sub-model for random masking, and the sub-model receives a training signal from the main-model similar to the self-distillation [68, 62, 40]. While the random masking technique amplifies the difficulty of the training process, this is counterbalanced by self-distillation loss since the outputs of the main-model are relaxed and easier objective than the ground-truth label. In summary, MaskSub applies a mask augmentation separated from the main-model, utilizing a relaxed loss form.
We analyze MaskSub using 100 epochs training on ImageNet [10]. Without MaskSub, loss convergence speed and corresponding accuracy are significantly degraded when mask augmentation is applied. Conversely, MaskSub successfully mitigates potential harmful effects from additional regularization, leading to a network training process that is even more efficient than standard training procedures. Also, MaskSub is not limited to mask augmentation and can be used for general drop regularizations. As a result, MaskSub is expanded to any random drop regularization without disrupting the convergence of original train loss; we employ three in-network drop-based options to show the applicability: masking [21, 1], dropout [44], and drop-path [27, 16]. Corresponding to each respective regularization strategy, we denote them MaskSub, DropSub, and PathSub. Among the three variants, MaskSub notably exhibits a significant enhancement in performance, which demonstrates necessity of mask augmentation in supervised learning.
We extensively validate the performance of MaskSub training. MaskSub is applied on various state-of-the-art supervised learning cases including DeiT-III [47], MAE finetuning [21], BEiTv2 finetuning [38], CLIP fine-tuning [14], ResNet-RSB [54], and Swin transformer [36]. MaskSub demonstrates remarkable performance improvement in all benchmarks. We argue that MaskSub can be regarded as a significant advancement as a novel way to utilize regularization for visual recognition.
@@
Figure 1: Overview of Masked Sub-model (MaskSub). (a) standard supervised training; (b) masking augmentation training. The masking is applied to the main model, which degrades performance; (c) our MaskSub training, which separates the masking from the main model using the sub-model and relaxes loss with self-distillation. MaskSub substantially improves the state-of-the-art training recipes [47, 54].
@@ 2 Related Work
Training recipe has been considered an important ingredient in building a high-performance network. He et al. [22] demonstrate that the training recipe significantly influences the network performance. RSB [54] is a representative and high-performance recipe for ResNet. With the emergence of ViT [15], the training recipe for ViT has gained the attention of the field. DeiT [45] shows that ViT can be trained to strong performance with only ImageNet-1k [10]. DeiT-III [47] is an improved version of DeiT, which applies findings from RSB to DeiT instead of distillation from CNN teacher. It is challenging to implement stronger or additional regularization in existing training recipes. To address this issue, we propose MaskSub approach employing sub-models.
@@
CoSub [48] introduces a similar concept to ours, utilizing sub-models. However, the objective of the sub-model differs significantly: while MaskSub aims to stabilize training through additional regularization, CoSub aims to train the sub-models in a collaborative manner [65]. We regard MaskSub as a more generalized framework since CoSub only considers the drop-path method to employ sub-models, whereas MaskSub can cover a variety of drop-based techniques including random masking.
@@
Self-distillation utilizes supervision from a network itself instead of using a teacher. ONE [68] uses a multi-branch ensemble to build superior output for the network and distill ensemble outputs as supervision for each branch. Some studies [62, 40] utilize the early-exit network for self-distillation. Those studies improve performance by using a full network as a teacher and an early-exit network as a student. MaskedKD [43] utilizes masking to reduce computation for knowledge distillation. From a self-distillation perspective, MaskSub presents a new insight to construct the student model (i.e., sub-model) from the teacher model (i.e., main-model) utilizing drop-based techniques. Note that most self-distillation studies are not compatible with recent training recipes [54, 47]. Thus, the general applicability of MaskSub is a significant contribution.
Self-supervised learning shares components with MaskSub. Previous works on contrastive learning incorporate two models with self-distillation loss [6, 18]. Want et al. [51] introduce a double tower with weak and strong augmentation for each model. In masked image models, supervised MAE [34] introduces additional supervised learning tasks to the MAE framework and accelerates MAE. Those studies partially share the fundamental concept with MaskSub and gave us the motivation. However, since training techniques applied to supervised learning are substantially different from techniques for self-supervised learning, MaskSub is significantly different from self-supervised learning approaches.
@@ 3 Method
We propose our method Masked Sub-model (MaskSub) with formulation and pseudo-code in Section 3.1. Section 3.2 presents analyses of MaskSub with loss convergence, accuracy, and gradient. In Section 3.3, we introduce variants of MaskSub: DropSub, and PathSub.
3.1 Masked Sub-model (MaskSub)
@@
Typically, a network with mask augmentation is trained with Eq. 2. But, we conjecture that training using Eq. 2 with high masking ratio (i.e. $r\geq 0.5$) may interfere with loss convergence and induce instability in training. To ensure training stability, we utilize the model output of equation Eq. 1, $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=0)$, as guidance for masking augmentation $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=r)$ instead of $\mathbf{y}_{i}$. In other words, Eq. 2 is changed as
@@
In our Masked Sub-model (MaskSub), the average of Eq. 1 and Eq. 3 is used as a loss function for the network. We designate $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=0)$ as the main-model and $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=r)$ as the sub-model. This naming convention is employed because a network with masked inputs appears to be a subset of the entire network. In Eq. 3, the main-model output $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=0)$ is used with stop-gradient. Thus, the sub-model is trained to mimic the main model, but we want the gradient for the main-model to be independent of the sub-model. This can be interpreted as self-distillation, where knowledge is transferred from the teacher (main-model) to the student (sub-model). Note that MaskSub can easily be expanded to binary cross-entropy loss by replacing the softmax function with the sigmoid function, which is used for recent training recipes [54, 47].
Algorithm 1 describes PyTorch-style pseudo-code of training with MaskSub. The masking ratio is put into the network input. The gradients are calculated on the average losses from the main and the sub-model. Note that MaskSub does not use any additional data augmentation, optimizer steps, and network parameters for the sub-model. In the implementation, we use MAE-style random masking [21], removing masked tokens to reduce computation costs, by default. It significantly reduces the training cost of MaskSub. Approximately, MaskSub with 50% masking requires $\times$1.5 computation to standard training. We will show the impact of MaskSub on diverse cases in Section 4, including computation analysis in Section A.1.
@@
    o1 = f(x)         # main model
    o2 = f(mask(x,r)) # sub-model (mask ratio: r)
@@
Figure 2: MaskSub training analysis. We use 50%-random masking to compare three training settings: standard Eq. 1, masking Eq. 2, and MaskSub. We visualize (a) accuracy on the validation set; (b) train loss without masking (standard); (c) train loss with masking.
In MaskSub training, the sub-model mimics the main-model, and it automatically controls the difficulty of sub-model. If the main-model produces output closely aligned with the ground-truth label, the sub-model loss aims to attain an accurate classification output under masking augmentation. Conversely, if the main-model fails to converge, the sub-model loss becomes considerably easier than constructing a ground-truth label. In summary, MaskSub prioritizes the learning process, ensuring that masking augmentation is exclusively applied to images that produce successful output in a standard setting. We assert that the prioritized loss mechanism of MaskSub enables the network to preserve its convergence speed and learning stability while maintaining the benefits of masking augmentation.
@@ 3.2 Analysis
We analyze MaskSub with training ViT-B [15] for 100 epochs on ImageNet-1k [10]. Based on DeiT-III [47] training, we shorten the epoch to 100 epochs and use image resolution $224\times 224$. We compare three settings: standard, masking, and MaskSub. The standard uses Eq. 1 as the training loss, and masking augmentation is not used. For the masking setting, the network is trained with Eq. 2. Note that it is a common practice to use a regularization or an augmentation in supervised learning. We compare those two settings with MaskSub. For analysis, we measured Eq. 1 ‘train loss - standard’ and Eq. 2 ‘train loss - masking’ regardless of loss used for training. It shows how losses changed by training setting.
Figure 3: Gradients magnitude of MaskSub. The figure shows gradient norms for the main- and sub-model of MaskSub. The gradient norm is the averaged value over all parameters for an epoch.
Figure 2 shows loss and accuracy trends in random masking 50% (i.e. $r_{mask}=0.5$) case. When random masking is applied to training (green), the masking loss (Figure 1(c)) converges better than the standard (blue). However, it significantly degrades the standard train loss (Figure 1(b)), resulting in a drop in accuracy (Figure 1(a)). Regularization over a certain strength often causes malicious effects on standard train loss, which decreases accuracy. As shown in Figure 1(b) and 1(c), MaskSub improves the loss convergence for both losses: original and masking, which brings a significant improvement in accuracy.
Figure 3 explains the learning pattern between main-model and sub-model of MaskSub (Eq. 3) in the aspect of gradients magnitude for training with random masking 50%. The gradient magnitude from the main-model (MaskSub-Main) is similar to that of other training. In contrast, gradients from the sub-model (MaskSub-Sub) have a small magnitude at the early stage. As the learning progresses, the gradients from the sub-model increase. It shows that MaskSub trains the network following our intention: automatic difficulty control. During the early stage of training, the gradients from the main-model lead the training. Following the progress of the main-model training, the sub-model adaptively increases its gradient magnitude and produces a reasonable amount of gradients at the end of training. In other words, the model training is relaxed from difficult masked inputs at the early stage, while it starts to learn masked input when the original inputs are sufficiently trained. We claim that the automatic difficulty control of MaskSub could be a general solution to introduce strong augmentation for supervised learning.

### v3

# Masking meets Supervision: A Strong Learning Alliance

@@ Abstract
Pre-training with random masked inputs has emerged as a novel trend in self-supervised training. However, supervised learning still faces a challenge in adopting masking augmentations, primarily due to unstable training. In this paper, we propose a novel way to involve masking augmentations dubbed Masked Sub-branch (MaskSub). MaskSub consists of the main-branch and sub-branch, the latter being a part of the former. The main-branch undergoes conventional training recipes, while the sub-branch merits intensive masking augmentations, during training. MaskSub tackles the challenge by mitigating adverse effects through a relaxed loss function similar to a self-distillation loss. Our analysis shows that MaskSub improves performance, with the training loss converging faster than in standard training, which suggests our method stabilizes the training process. We further validate MaskSub across diverse training scenarios and models, including DeiT-III training, MAE finetuning, CLIP finetuning, BERT training, and hierarchical architectures (ResNet and Swin Transformer). Our results show that MaskSub consistently achieves impressive performance gains across all the cases. MaskSub provides a practical and effective solution for introducing additional regularization under various training recipes. Code available at https://github.com/naver-ai/augsub
@@ 1 Introduction
Supervised learning is the most basic and effective way to train a network to achieve high performance on a target task. To improve supervised learning, diverse regularizations are developed and used as training recipes [48, 50, 57], which represent a group of sophisticatedly tuned regularizations to maximize learning performance. Supervised learning has always held an advantage over self-supervised learning [7, 4] based on the benefit of supervision. However, emergence of Vision Transformer (ViT) [16] and Masked Image Modeling (MIM) [1, 22, 41] is changing this trends. ViT, which lacks inductive bias compared to convolution networks, poses many challenges to generalization performance for supervised learning. On the other hand, MIMs such as MAE [22] rise as an alternative pretraining method for ViT by achieving competitive performance with supervised learning recipes. Although a recent study [50] shows that new supervised learning outperforms MIMs, the gap is insignificant. Thus, MIMs are still a strong competitor of supervised learning methods.
@@
MIM masks random areas of an input image and forces the network to infer the masked area using the remaining area. A representative part of MIM is high mask ratios over 50%. Although MIM also works at small mask ratios, it shows remarkable performance when trained with a high mask ratio. The high mask ratio is a major difference between MIM and supervised learning since this high mask ratio, over 50%, is not beneficial in supervised learning. Supervised learning also has utilized random masking as an augmentation [18, 70], but it significantly degrades performance when the masking ratio is high. In other words, supervised learning is not applicable for strong masking augmentation. We conjecture that it is a major problem of the current supervised learning recipe, and there is room for improvement by enabling strong masking.
@@
Figure 1: Overview of Masked Sub-branch (MaskSub). (a) standard supervised training; (b) masking augmentation training. The masking is applied to the main model, which degrades performance; (c) our MaskSub training, which separates the masking from the main model using the sub-branch and relaxes loss with self-distillation. MaskSub substantially improves the state-of-the-art training recipes [50, 57].
Our goal is to improve supervised learning with strong mask augmentation over 50%. To this end, we introduce a novel learning framework using a “sub-branch” alongside the main-branch; throughout this paper, we use the term “sub-branch” to describe a model with dropped inputs. The main-branch uses standard training recipes [57, 50], while the sub-branch utilizes mask augmentation. We name our method as Masked Sub-branch (MaskSub).
We visualize the overview of MaskSub in Figure 1. We consider a high masking ratio over 50% as similar in MAE [22]. Figure 1 (b) shows that applying the strong random masking on the main-branch may lead to degraded performance. In contrast, as in Figure 1 (c), MaskSub leverages the sub-branch for random masking, and the sub-branch receives the training signal from the main-branch similar to the self-distillation [72, 66, 43]. While the random masking technique amplifies the difficulty of the training process, this is counterbalanced by self-distillation loss since the outputs of the main-branch are relaxed and easy-to-learn objective than the ground-truth label. In summary, MaskSub applies a mask augmentation separated from the main-branch, utilizing a relaxed loss form.
We analyze MaskSub using 100 epochs training on ImageNet [11]. Without MaskSub, loss convergence speed and corresponding accuracy are significantly degraded when mask augmentation is applied. Conversely, MaskSub mitigates potential harmful effects from additional regularization, leading to a network training process that is even more efficient than standard training procedures. Also, MaskSub is not limited to mask augmentation and can be used for general drop regularizations. As a result, MaskSub is expanded to any random drop regularization without disrupting the convergence of original train loss; we employ three in-network drop-based options to show the applicability: masking [22, 1], dropout [47], and drop-path [28, 17]. Corresponding to each respective regularization strategy, we denote them MaskSub, DropSub, and PathSub. Among the three variants, MaskSub notably exhibits a remarkable performance enhancement, demonstrating the necessity of mask augmentation in supervised learning.
We extensively validate the performance of MaskSub. MaskSub is applied on various state-of-the-art supervised learnings including DeiT-III training [50], MAE finetuning [22], BEiTv2 finetuning [41], CLIP finetuning [15], BERT training [13], ResNet-RSB [57], and Swin transformer [38]. MaskSub demonstrates remarkable performance improvement in all benchmarks. We argue that MaskSub can be regarded as a novel way to utilize regularization for visual recognition.
@@ 2 Related Work
Training recipe has been considered an important ingredient in building a high-performance network. He et al. [23] demonstrate that the training recipe significantly influences the network performance. RSB [57] is a representative and high-performance recipe for ResNet. With the emergence of ViT [16], the training recipe for ViT has gained the attention of the field. DeiT [48] shows that ViT can be trained to display strong performance with only ImageNet-1k [11]. DeiT-III [50] is an improved version of DeiT, which applies findings from RSB to DeiT instead of distillation from CNN teacher. It is challenging to implement stronger or additional regularization in existing training recipes. To address this issue, we propose our MaskSub employing sub-branchs.
@@
CoSub [51] introduces a similar concept to ours, utilizing sub-branchs. However, the sub-branch objective differs: while MaskSub aims to stabilize training through additional regularization, CoSub aims to train the sub-branchs by co-training [69]. We regard MaskSub as a more generalized framework since CoSub only considers the drop-path method to employ sub-branchs, whereas MaskSub can cover various drop-based techniques, including masking.
@@
Self-distillation utilizes supervision from a network itself instead of using a teacher. ONE [72] uses a multi-branch ensemble to build superior output for the network and distill ensemble outputs as supervision for each branch. Some studies [66, 43] utilize the early-exit network for self-distillation. Those studies improve performance by using an entire network as a teacher and an early exit network as a student. MaskedKD [46] utilizes masking to reduce computation for knowledge distillation. From a self-distillation perspective, MaskSub presents a new insight into constructing the student model (i.e., sub-branch) from the teacher model (i.e., main-branch) utilizing drop-based techniques. Note that most self-distillation studies are not compatible with recent training recipes [57, 50]. Thus, the general applicability of MaskSub is a notable contribution.
Self- and semi-supervised learning share components with MaskSub. Contrastive learning incorporates two models with self-distillation loss [6, 19]. Want et al. [54] introduce a double tower with weak and strong augmentation for each model. MAE [22] uses masked image reconstruction as self-supervision, and supervised MAE [36] introduces supervised learning as an additional task for MAE. MAE and supMAE aim to reconstruct masked images using MAE training recipe, rather than supervised learning. In contrast, MaskSub only relies on label-related loss with a supervised learning recipe. In semi-supervised learning, UDA [61] introduces a two-branch framework, similar to the main- and sub-branch in MaskSub. However, MaskSub is more computationally efficient by using masking [22] and removing label-consistency checks for unlabeled data. Also, MaskSub extends the two-branch framework to supervised learning via distillation loss, in contrast to UDA’s consistency loss. While these studies share the fundamental concept with MaskSub and inspired our work, the training techniques for supervised learning differ from those in semi- and self-supervised learning. Thus, we argue that MaskSub retains its originality and novelty compared to these studies.
@@ 3 Method
We propose our method Masked Sub-branch (MaskSub) with formulation and pseudo-code in Section 3.1. Section 3.2 presents analyses of MaskSub with loss convergence, accuracy, and gradient. In Section 3.3, we introduce variants of MaskSub: DropSub, and PathSub.
3.1 Masked Sub-branch (MaskSub)
@@
Typically, a network with mask augmentation is trained with Eq. 2. But, we conjecture that training using Eq. 2 with a high masking ratio (i.e. $r\geq 0.5$) may interfere with loss convergence and induce instability in training. To ensure training stability, we utilize the model output of equation Eq. 1, $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=0)$, as guidance for masking augmentation $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=r)$ instead of $\mathbf{y}_{i}$. In other words, Eq. 2 is changed as
@@
In our Masked Sub-branch (MaskSub), the average of Eq. 1 and Eq. 3 is used as a loss function for the network. We designate $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=0)$ as the main-branch and $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=r)$ as the sub-branch. This naming convention is employed because a network with masked inputs appears to be a subset of the entire network. In Eq. 3, the main-branch output $f_{\theta}(\mathbf{x}_{i}|r_{\mathrm{mask}}=0)$ is used with stop-gradient. Thus, the sub-branch is trained to mimic the main model, but we want the gradient for the main-branch to be independent of the sub-branch. This can be interpreted as self-distillation, where knowledge is transferred from the teacher (main-branch) to the student (sub-branch). Note that MaskSub can easily be expanded to binary cross-entropy loss by replacing the softmax function with the sigmoid function, which is used for recent training recipes [57, 50].
Algorithm 1 describes PyTorch-style pseudo-code of training with MaskSub. The masking ratio is put into the network input. The gradients are calculated on the main and sub-branch average losses. Note that MaskSub does not use any additional data augmentation, optimizer steps, and network parameters for the sub-branch. We use MAE-style random masking [22], removing masked tokens to reduce computation costs by default. It significantly reduces the training cost of MaskSub. Approximately, MaskSub with 50% masking requires $\times$1.5 computation to standard training. In practical implementation, we separate the main and sub-branch backward passes utilizing the gradient accumulation of PyTorch. So, VRAM for the main-branch can be released before the sub-branch computation, which eliminates the need for additional VRAM for MaskSub training. We will show the impact of MaskSub on diverse cases in Section 4, including computation analysis in Section A.2.
@@
o1 = f(x) # main
o2 = f(mask(x,r)) # sub (mask ratio: r)
@@
(d) Gradient norm
Figure 2: MaskSub training analysis. We use 50%-random masking to compare three training settings: standard Eq. 1, masking Eq. 2, and MaskSub. We visualize (a) validation accuracy; (b) train loss without masking; (c) train loss with masking; (d) gradient norm.
MaskSub automatically controls the difficulty of the sub-branch. If the main-branch is close to the ground-truth label, the sub-branch loss aims to attain the ground-truth label under masking. Conversely, if the main-branch fails to converge, the sub-branch loss becomes easy. This difficulty design is inspired by distillation studies [39, 8, 29]. The distillation becomes difficult when a high-performance network is used as a teacher [39, 8]. An early-stage network is easy, and an end-stage network is challenging [29]. Thus, MaskSub can be considered as a sample-wise masking augmentation that is exclusively applied to images that produce successful output in the main-branch.
@@ 3.2 Analysis
We analyze MaskSub with ViT-B [16] for 100 epochs training on ImageNet-1k [11]. Based on DeiT-III [50], we shorten the epoch to 100 epochs and use image resolution $224\times 224$. We compare three settings: standard, masking, and MaskSub. The standard uses Eq. 1 as the training loss, and masking augmentation is not used. For the masking setting, the network is trained with Eq. 2. Note that it is a common practice to use a regularization or an augmentation in supervised learning. We compare those two settings with MaskSub. For analysis, we measured Eq. 1 ‘train loss - standard’ and Eq. 2 ‘train loss - masking’. It shows how losses changed by training setting.
Figure 2 shows loss and accuracy trends in random masking 50% (i.e., $r_{mask}=0.5$) case. When random masking is applied to training (green), the masking loss (Figure 2(c)) converges better than the standard (blue). However, it significantly degrades the standard train loss (Figure 2(b)), resulting in a drop in accuracy (Figure 2(a)). Regularization over a certain strength often causes malicious effects on standard train loss, which decreases accuracy. As shown in Figure 2(b) and 2(c), MaskSub improves the loss convergence for both losses, original and masking, which brings an improvement in accuracy.
Figure 2(d) explains the learning pattern between main-branch and sub-branch of MaskSub (Eq. 3) in the aspect of gradients magnitude for training with random masking 50%. The gradient magnitude from the main-branch (MaskSub-Main) is similar to that of other training. In contrast, gradients from the sub-branch (MaskSub-Sub) have a small magnitude at the early stage. As the learning progresses, the gradients from the sub-branch increase. It shows that MaskSub trains the network following our intention: automatic difficulty control. During the early stage of training, the gradients from the main-branch lead the training. Following the progress of the main-branch training, the sub-branch adaptively increases its gradient magnitude and produces a reasonable amount of gradients at the end of training. In other words, the model training is relaxed from challenging masked inputs at the early stage, while it starts to learn masked input when the original inputs are sufficiently trained. We claim that the automatic difficulty control of MaskSub could be a general solution to introduce strong augmentation for supervised learning.